Write the equation of the line that passes through point (6,4) with a slope of 2?
Answer by fractalier(6550)

Use the point-slope form
y - y1 = m(x - x1)
y - 4 = 2(x - 6)
y - 4 = 2x - 12
y = 2x - 8